OpenStudy (anonymous):

cos(2theta+pi/3)=-radical(3)/2

OpenStudy (anonymous):

\[ \cos\left(2\theta+{\pi\over3}\right)={\sqrt{3}\over2} \] this happens when \[2\theta+\frac{\pi}{3}={\pi\over6}+n\pi\\ 2\theta=n\pi-{\pi\over6}\\ \theta=n{\pi\over2}-{\pi\over12} \]

OpenStudy (anonymous):

if you have the domain of theta, use the appropriate value for "n"